What is the tangential acceleration of the object if it accelerates uniformly in circular motion with changes in velocity of 80 m/s and in the total changes of time of 25 seconds?
1
Expert's answer
2020-12-03T05:16:10-0500

"a_\tau=\frac{\Delta v}{\Delta t}=\frac{80}{25}=3.2(m/s^2)"